• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name cusanno

    Log entry time 08:02:01 on March18,2006

    Entry number 167265

    keyword=Owl Shift Summary

    Shift started just after production was restored. Something bad in the beam charge asymmetry. Bob Michaels came in to look at it.
    Furthermore there are still some problems with T3 efficiency, due to timing as far as I understand.

    00:50 - Qasy feedback restarted. Beam charge asymmetry now looks fine.

    01:11 - testing T3 efficiency

    01:35 - switch back to standard production

    02:10 - NMR measurement performed

    03:10 - some minor problems with DAQ. One-run test at 0.5 uA, ps2=1000000, then switch back to production, ps2=60.

    04:51 - CODA problems, same as before: end-of-run failed, start-run not working. ROCs and CODA components all rebooted.
    We take chance to do NMR measurement, scheduled at 5.

    After Bob "repair", beam charge asymmetry was always in the specifications (<50 ppm), as well as SLI, always around 8 x 10-5


    Production run:

    3100, 1M events, undestanding Beam Charge Asymmetry, first run after switching in (temptative) production
    3101, 1M events
    3102, 930k events, Qasy feedback rebooted at the end
    3103, 930k events, Beam Charge Asymmetry OK, but still checking T3
    3104, 1M events, ps1=100000, ps2=1000, ps3=1, T3 check
    3105, 1M events, ps1=100000, ps2=60, ps3=1, standard production
    3106, 1M events
    3107, 1M events
    3108, 450k events, ended to to test al low current
    3109 JUNK
    3110, 50 k events, test at 0.5 uA, ps2= one million
    3111, 1 M events, back to production
    3112, 1 M events
    3113, 1 M events
    3114, 160 k events, ended to check raster size
    3115, 1M events, end-of-run failed
    3116 JUNK
    3117, 1M events
    3118, 1M events
    3119, 1M events
    3120, 1M events
    3121, 1M events
    3122, 1M events
    3123, 1M events
    3124, 1M events
    3125, 1M events